OpenStudy (anonymous):

I do not understand the concepts of Arithmetic sequences. Please explain if you can how to do this type of math.

OpenStudy (turingtest):

the *concept* is that the difference between any two successive terms is the same (called the "common difference")

OpenStudy (turingtest):

for example: 1,3,5,7,9,11 is arithmetic because 11-9=2 9-7=2 7-5=2 3-1=2 the common difference d of this arithmetic sequence (AS) is 2

OpenStudy (turingtest):

other example: 4,9,14,19,24,29,... the ... part means it continues in that way this sequence is arithmetic what is the common difference between each successive term?

OpenStudy (anonymous):

which sequence..........ap ,gp???

OpenStudy (anonymous):

atleast tell me which sequence?????????

OpenStudy (anonymous):

is 5 the common difference between the numbers

OpenStudy (turingtest):

she wrote arithmetic @Mayrie yes

OpenStudy (turingtest):

good job :)

OpenStudy (anonymous):

I do not understand how the problems are calculated using formulas for example: Find the eighth term of an arithmetic sequence that begins with 5 and has a common difference of 9. Substitute in the formula using a1 = 5, n = 8, and d = 9. an = a1 + (n − 1)d a8 = 5 + (8 − 1)(9) = 5 + (7)(9) = 5 + 63 = 68 How is this answer gotten. I am looking at it but it does not make any sense to me.